OpenStudy (anonymous):

Input the equation of the given line in standard form. The line through (1, -4) and parallel to 2x + 3y =4.

OpenStudy (anonymous):

It should be like \[ 2x + 3y =c \] Use x= and y=-4 to get c

OpenStudy (anonymous):

x=1 and y=-4 gives 2(1) + 3(-4) =c 2 -12=c c=-10 The equation is 2x + 3y =-10